7. Find the inverse Laplace Transform of X(so2 with ROC-1< Rels) 1.
Answer #1

Given the Laplace transform is

X(s)=\frac{-5s-7}{(s+1)(s-1)(s+2)}

Let's split the above expression into partial fractions.

\frac{-5s-7}{(s+1)(s-1)(s+2)}=\frac{A}{s+1}+\frac{B}{s-1}+\frac{C}{s+2}

\Rightarrow -5s-7=A(s-1)(s+2)+B(s+1)(s+2)+C(s+1)(s-1)

\Rightarrow -5s-7=A(s^2+s-2)+B(s^2+3s+2)+C(s^2-1)

\Rightarrow -5s-7=s^2(A+B+C)+s(A+3B)+(-2A+2B-C)

Comparing the coefficients on both sides

\\A+B+C=0 \\A+3B=-5 \\-2A+2B-C=-7

Solving the above set of three equations

A=1;\ B=-2\ \&\ C=1

So,

X(s)=\frac{1}{s+1}+\frac{-2}{s-1}+\frac{1}{s+2}=X_1(s)+X_2(s)+X_3(s)

Laplace transform, that is expressed in partial fractions as X(s)=\sum_{j=1}^m\frac{A_j}{s+a_j} ,

For each \frac{A_j}{s+a_j} ,

if the ROC is right to the pole at s=-a_i, the inverse Laplace transform is  A_ie^{-a_it}u(t)

if the ROC is left to the pole at s=-a_i , the inverse Laplace transform is -A_ie^{-a_it}u(-t)

In the given problem, we have the region of convergence (ROC) at -1<\text{Re}\left \{ s \right \}<1

For

X_1(s)=\frac{1}{s+1}, the pole is at s=-1.

So, the ROC is to the right of the poles. Hence the inverse Laplace transform is

x_1(t)=e^{-t}u(t)

For

X_2(s)=\frac{-2}{s-1}, the pole is at s=1 .

So, the ROC is to the left of the poles. Hence the inverse Laplace transform is

x_2(t)=2e^{t}u(-t)

For

X_3(s)=\frac{1}{s+2}, the pole is at s=-2 .

So, the ROC is to the right of the poles. Hence the inverse Laplace transform is

x_3(t)=e^{-2t}u(t)

So, the inverse Laplace transform of X(s) is

x(t)=x_1(t)+x_2(t)+x_3(t)

\Rightarrow x(t)=e^{-t}u(t)+2e^{t}u(-t)+e^{-2t}u(t)

Hence, the inverse Laplace transform of X(s)=\frac{-5s-7}{(s+1)(s-1)(s+2)} with region of convergence (ROC) -1<\text{Re}\left \{ s \right \}<1 is

x(t)=e^{-t}u(t)+2e^{t}u(-t)+e^{-2t}u(t)